The student council is hosting a drawing to raise money for scholarships. They are selling tickets for $1 each and will sell 1500 tickets. There is one $600 grand prize, five $50 second prizes, and twenty $25 third prizes. You just bought a ticket. Find the expected value for your profit. Round your answer to the nearest cent.
Answer by Theo(13342)

the probability that the ticket will cost you 1 dollar is 1.

the probability that the ticket will win you 600 dollars * 1/1500 = 600/1500.

the probability that the ticket will win you 50 dollars is 50 * 5/1500 = 250/1500.

the probability that the ticket will win you 25 dollars is 25 * 20/1500 = 500/1500.

your expected value is -1 + 600/1500 + 250/1500 + 500/1500 = -.1.

the average loss for each dollar that is bet is equal to 10 cents.